Homework: Assignment 11 Score: 0 of 1 pt 8.5.4 1 of 25 (17 complete) Score: 61.67%, Ques For the experiment described below, let x determine a random variable, and use your knowledge of probability to prepare a probability distribution. Three balls are drawn (with replacement) from a bag that contains 5 orange balls and 3 green balls. The number of green balls is counted. A probability distribution for the given experiment is as follows. (Round to four decimal places as needed.) P(x)
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

8.5.4

Prob that you get a green ball is 3 out of 8, as there are 3 green balls out of 8 balls.

P(x=0) = 3C0*(3/8)^0 *(5/8)^3 = .2441

P(x=1) = 3C1*(3/8)^1 *(5/8)^2 = 0.4395

P(x=2) = 3C2*(3/8)^2 *(5/8)^1 = 0.2637

P(x=3) = 3C3*(3/8)^3 *(5/8)^0 = 0.0527